THE PHOTODISSOCIATION DYNAMICS OF METHYL NITRATE

\underline{CRAIG MURRAY}, ERIKA L. DERRO, MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104; and MARK D. MARSHALL, Department of Chemistry, Amherst College, Amherst, MA 01002.

Interest in the photodissociation mechanism of alkyl nitrates (RONO_{2}, where R = CH_{3}, C_{2}H_{5}, etc.) stems from recent experimental measurements in the troposphere indicating that they are an important component of “missing NO_{y}”. In this study, the photodissociation dynamics of methyl nitrate, CH_{3}ONO_{2}, at 193 nm have been investigated by examining the products from the primary channel, namely, CH_{3}O and NO_{2}. Laser-induced fluorescence (LIF) spectroscopy was employed to probe the nascent internal energy distribution of the CH_{3}O radical, a small fraction of which was found to be produced with one quantum of C−O stretch excitation. The stretch-excited methoxy was observed to be formed with a significantly greater degree of rotational excitation than the vibrational ground state. Furthermore, dispersed fluorescence measurements reveal that the NO_{2} fragment is produced electronically excited with internal energies out to the NO + O dissociation limit, indicating that the initial excitation is strongly localized on the NO_{2} moiety. Comparisons will be drawn with the analogous photodissociation of nitric acid, HONO_{2}.